OpenStudy (gazela1997):

MATH EXAM TOMORROW HELP MEDAL AND FAN Tickets for a popular rock concert where sold out for an auditorium that has 8000 seats. The premium seat tickets sell for 50 and regular seats are 30. If the total revenue generated from sales was 250,00, determine the number of tickets were sold at each price.

OpenStudy (gazela1997):

need 2 linear equations to solve this problem

OpenStudy (anonymous):

Let p be the number of premium seat tickets sold, and r be the number of regular seat tickets sold. What two equations can we set up?

OpenStudy (gazela1997):

idk

OpenStudy (anonymous):

Okay, no worries. So we know 8000 total tickets were sold, right?

OpenStudy (gazela1997):

that why Im asking you lol

OpenStudy (gazela1997):

can you just show me how you do it please?

OpenStudy (anonymous):

Yeah I can explain it. So we know 8000 total tickets were sold, so that means the number of premium seat tickets and regular seat tickets combined have to be 8000, right?

OpenStudy (gazela1997):

k

OpenStudy (anonymous):

So, since we let p be the number of premium seat tickets, and r be the number of regular seat tickets, we can set up the first equation \[p + r = 8000\]

OpenStudy (gazela1997):

ok

OpenStudy (anonymous):

Now, we look at the revenue gained. The total revenue gained was 250,000. So we know that this total comes from the price of each ticket multiplied by their respective price combined. Since the premium seat tickets cost 50 and the regular seat tickets cost 30, we can set up the second equation \[50p + 30r = 250,000\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

I made a mistake above in what I said. I meant to say that the total revenue gained comes from the price of each ticket multiplied by the number of those tickets all combined.

OpenStudy (gazela1997):

it makes sesne but how do i solve the problem with these equations?

OpenStudy (anonymous):

Okay so we have the equations \[p + r = 8000\] and \[50p + 30r = 250,000\] now we can take the first equation and rearrange the variables to get \[p = 8000-r\] right?

OpenStudy (gazela1997):

just do it please

OpenStudy (anonymous):

I am, I'm just trying to walk you through it as well. So now that we have an equation that is solved for p, we can substitute it into the second equation to get \[50[8000-r] + 30r = 250,000\] and then just solve for r, and you'll get a number that you can put into the first equation and solve for p, and you'll have the answer.

OpenStudy (gazela1997):

where di p go?

OpenStudy (gazela1997):

did*

OpenStudy (anonymous):

I used the rearranged first equation \[p = 8000-r\] and substituted it for p in the second equation \[50p + 30r = 250,000\]

OpenStudy (gazela1997):

got it

OpenStudy (anonymous):

Does it all make sense?